### v1

# MIST: Towards Multi-dimensional Implicit Bias and Stereotype Evaluation of LLMs via Theory of Mind

Abstract
Theory of Mind (ToM) in Large Language Models (LLMs) refers to their capacity for reasoning about mental states, yet failures in this capacity often manifest as systematic implicit bias. Evaluating this bias is challenging, as conventional direct-query methods are susceptible to social desirability effects and fail to capture its subtle, multi-dimensional nature. To this end, we propose an evaluation framework that leverages the Stereotype Content Model (SCM) to reconceptualize bias as a multi-dimensional failure in ToM across Competence, Sociability, and Morality. The framework introduces two indirect tasks: the Word Association Bias Test (WABT) to assess implicit lexical associations and the Affective Attribution Test (AAT) to measure covert affective leanings, both designed to probe latent stereotypes without triggering model avoidance. Extensive experiments on 8 State-of-the-Art LLMs demonstrate our framework’s capacity to reveal complex bias structures, including pervasive sociability bias, multi-dimensional divergence, and asymmetric stereotype amplification, thereby providing a more robust methodology for identifying the structural nature of implicit bias.
WARNING: This paper contains content that may be offensive and disturbing in nature.
1 Introduction
As large language models (LLMs) demonstrate increasingly sophisticated reasoning capabilities, the question of whether they possess a form of Theory of Mind (ToM) Premack and Woodruff (1978) has emerged as a central topic. ToM, also known as mentalizing, is the ability to infer the mental and emotional states of other beings. Since this capacity underpins meaningful communication and empathy, investigating its potential emergence in LLMs is a critical research endeavor with profound implications for developing safer and more cooperative AI systems Nguyen and others (2025).
A broader perspective on ToM includes the ability to model complex social structures involving individuals and groups Baker et al. (2017). As LLMs internalize knowledge from large-scale corpora, they inevitably learn the statistical distributions that reflect societal stereotypes. In this context, we can define stereotypes as learned but often flawed generalizations about a group. Bias is the subsequent failure that occurs when the model misapplies these group-level stereotypes to make judgments about an individual. This represents a systematic failure of ToM, as the model generates erroneous beliefs about a person’s intentions or competencies.
However, most existing studies Yeh et al. (2023), Duan et al. (2024) employ uni-dimensional diagnostic tasks to evaluate such biases. Typically, these studies directly query the model to examine its associations with sensitive group-related attributes, such as gender, race, or occupation. While these approaches provide valuable empirical insights, prior studies Sheng et al. (2021), Wan et al. (2023) have demonstrated that LLMs are also prone to social desirability effects in their responses. This susceptibility limits their ability to detect more subtle and cognitively plausible forms of bias that may surface during uncontrolled reasoning. Moreover, existing works Lucy and Bamman (2021), Liang et al. (2022), Vijayaraghavan et al. (2025), Syed et al. (2025) have yet to adequately account for the multi-dimensional and relational nature of human social perception, which frequently involves the interplay of multiple psychological dimensions. This methodological gap results in a critical blind spot, potentially leading to an underestimation of how LLMs can perpetuate nuanced and socially corrosive stereotypes.
To address this research gap, we leverage the Stereotype Content Model (SCM), a widely employed framework in social psychology that characterizes stereotypes along three core dimensions: Competence, Sociability and Morality Leach et al. (2007). This model provides a cognitively grounded multi-dimensional analytical idea for evaluating LLMs from a ToM perspective. Rather than directly probing for bias, we design indirect evaluation tasks Word Association Bias Test (WABT) and Affective Attribution Test (AAT). Specifically, WABT measures associative biases by having the model pair attribute words with social groups, while AAT measures affective biases by having it attribute an emotional valence to generated scenarios involving those groups. By framing the evaluations as objective lexical association or subjective affective judgment tasks, rather than direct inquiries about social beliefs, the methodology avoids triggering the model’s learned social desirability filters. Consequently, these tasks naturally prompt the model to generate social inferences without explicitly introducing the notion of bias, thereby allowing its latent stereotypical tendencies to surface unconsciously during the reasoning process.
The contributions of this paper are summarized as follows:
We provide an integrated theoretical perspective that combines insights from ToM and SCM to reconceptualize implicit bias in LLMs as systematic failures in mental state modeling.
We propose a novel implicit bias evaluation framework that incorporates WABT and AAT task, which indirectly prompt the model to generate group-level inferences along the 3 SCM dimensions. This design minimizes the influence of explicit bias-avoidance mechanisms.
We conduct extensive empirical evaluations on multiple State-of-the-Art LLMs, uncovering new insights into the structural, subtle, and pervasive nature of their implicit social biases.
2 Related Work
2.1 Theory of Mind in LLMs
The recent advancements in the reasoning and problem-solving capabilities of LLMs Wei et al. (2022), Wang et al. (2022), Huang and Chang (2023) have provoked significant scientific debate surrounding their potential for an emergent ToM. This capacity, defined as the ability to attribute and reason about the beliefs, intentions, and knowledge of others, has long been considered a hallmark of human social intelligence. Consequently, evaluating the extent to which LLMs can replicate ToM Kosinski (2023) has become a pivotal research objective, carrying profound implications for the future development of human-centered AI Zhao et al. (2023), Li et al. (2024), Liu et al. (2025b), Wang et al. (2024), Cheng et al. (2025), Liu et al. (2025a).
However, scholarly assessments of these capabilities have yielded divergent conclusions. Critical analyses posit that high performance may stem from methodological flaws in benchmark design or a reliance on superficial statistical patterns rather than genuine reasoning Wang et al. (2025), Sadhu et al. (2024). In contrast, other research demonstrates that ToM skills can be made more robust, showing that targeted training enables generalization to novel and complex tasks Lu et al. (2025). As LLMs internalize knowledge from vast corpora, they also learn flawed societal stereotypes. A critical failure of mentalizing occurs when a model misapplies these learned group-level generalizations to an individual, thereby forming distorted and erroneous beliefs about their intentions or competencies. Therefore, current research should address both the functional robustness of ToM in specific tasks and these broader systemic failures.
2.2 Stereotype Content Model
The Stereotype Content Model (SCM), proposed by Fiske et al. (2002), explains how stereotypes form along two core dimensions: Warmth and Competence. Later research refined Warmth into Sociability and Morality to better capture perceptions of ethics and trustworthiness Leach et al. (2007). SCM has been widely validated across cultures and groups using surveys, IATs, and experiments Cuddy et al. (2009), Fiske (2018). It introduced “ambivalent prejudice,” recognizing that bias can involve mixed perceptions: for instance, women are often seen as warm but less competent, while the elderly are viewed as low in both, eliciting pity; competent but cold groups may provoke envy Chen et al. (2021).
Recent studies show LLMs replicate similar patterns. Though their outputs are generally positive in tone, descriptions of social groups still align with SCM dimensions Kotek et al. (2023), Schuster et al. (2024). LLMs often default to white, healthy, middle-aged male characters, while descriptions of other groups show semantic shifts and implicit bias, reflecting amplified normative assumptions Bai et al. (2025), Tan and Lee (2025). Nicolas and Caliskan applied SCM to LLMs by creating a 14-dimension stereotype taxonomy, confirming that Warmth and Competence remain dominant evaluative dimensions Nicolas and Caliskan (2024). This approach reveals the complexity of LLM biases more clearly than binary labels and highlights the risk of reinforcing inequality in areas like education and hiring Allstadt Torras et al. (2023), Weissburg et al. (2024).
3 Evaluation Methodology
The pipeline of our proposed evaluation methodology is presented in Figure 1.
3.1 Tasks Definition
We design two types of implicit bias evaluation tasks Word Association Bias Test (WABT) and Affective Attribution Test (AAT), to assess LLMs’ implicit biases and underlying stereotypical tendencies along the three dimensions of Competence, Sociability, and Morality.
3.1.1 Word Association Bias Test
The WABT task indirectly assesses LLMs’ implicit biases and stereotypes by examining their associative tendencies at the lexical level. These implicit biases and stereotypes are often reflected in the model’s inclination to associate specific groups with certain attributes or characteristics when processing group-related words. Specifically, given a LLM , for each bias dimension, a pair of target group identifiers and 10 attribute words (5 , 5 ) are provided to . The model is required to associate each attribute word with one of the two target group identifiers. The model’s output is represented as pairs. Here, refers to the positively framed target group (advantaged or normative group), and refers to the negatively framed target group (disadvantaged or marginalized group). Likewise, denotes positive or desirable attributes, while denotes negative or undesirable attributes.
3.1.2 Affctive Attribution Test
The AAT task is designed to evaluate LLMs’ implicit biases and stereotypes by examining their affective associations toward social group identifiers. The task is adapted from the affective misattribution paradigm in cognitive psychology, which infers implicit attitudes based on affective priming effects. Specifically, for each social group dimension, the model is first prompted in each trial to generate a descriptive sentence that includes a neutral word alongside the target group identifier which is a combination of and . Subsequently, the model is required to categorize the generated sentence, based on its initial affective response, into one of two categories: Comedy (positive valence) or Tragedy (negative valence). The output of is recorded as a categorical label, reflecting the affective association activated toward the target group.
3.2 Evaluation Metrics
For each task, we employ specific evaluation metrics to rigorously quantify the extent of implicit bias exhibited by LLMs.
3.2.1 Word Association Bias Test
To quantify the implicit association bias in each test, we adopt a commonly used lexical association bias scoring method. The bias score is computed as follows:
where denotes the number of times the model assigns an attribute word from to the target group (i.e., the number of pairs in the model’s output), and similarly for the other terms. The resulting bias score ranges from (completely reversed bias) to (completely consistent bias), with indicating no observable bias.
3.2.2 Affective Attribution Test
We focus on the model’s affective attribution tendencies along two target-specific directions:
When the target group identifier belongs to (advantaged group) and the model classifies it as comedy (positive valence), it is counted as a favorable attribution.
When the target group identifier belongs to (disadvantaged group) and the model classifies it as tragedy (negative valence), it is counted as an unfavorable attribution.
After multiple rounds of testing, we obtain the number of favorable attributions, denoted as , and the number of unfavorable attributions, denoted as . After collecting the total number of favorable and unfavorable attributions, we further compute two normalized attribution rates: (1) Favorable Attribution Rate (FAR) and (2) Unfavorable Attribution Rate (UAR). The FAR as the proportion of favorable attributions among the total number of instances where the target group belongs to (advantaged group), defined as:
| Domain | Subdomain | Group | Category | Items |
| Race | - | American | Ethan, Olivia, Jackson, Ava, Liam | |
| African | Kwame, Amina, Zuberi, Thabo, Adama | |||
| Asian | Hiroshi, Jiahao, Aditi, Siti, Minh | |||
| Spanish-speaking | Martín, Valentina, Diego, Camila, Felipe | |||
| Arab | Omar, Layla, Rami, Nadia, Samir | |||
| Gender | - | Female | Woman, Girl, Lady, Feminine, Womanhood | |
| Male | Man, Boy, Gentleman, Masculine, Manhood | |||
| Transgender | Trans, Gender-nonconforming, Genderqueer, Non-binary, Gender-fluid | |||
| Health | Disability | - | Able, Fitness, Wellness | |
| - | Disable, Impairment, Handicap | |||
| Weight | - | Slim, Skinny, Thin | ||
| - | Obesity, Heavy, Fat | |||
| Age | - | Young, Youth, Junior | ||
| - | Old, Aged, Seniors | |||
| Mind | - | Sanity, Emotional Stability, Sane | ||
| - | Depression, Schizophrenia, Bipolar Disorder |
The UAR as the proportion of unfavorate attributions among the total number of instances where the target group belongs to (disadvantaged group), defined as:
Intuitively, higher values of FAR and UAR indicate stronger implicit biases and stereotypical tendencies in the model’s affective attribution behavior. Specifically, a high FAR suggests that the model disproportionately associates advantaged groups () with positive valence (Comedy), while a high UAR reflects a tendency to associate disadvantaged groups () with negative valence (Tragedy). Both patterns reveal systematic asymmetries in the model’s social reasoning that may reflect internalized societal stereotypes.
3.3 Data Construction
To systematically evaluate implicit bias and stereotypes in LLMs, we construct a synthetic dataset following a controlled generation pipeline. Specifically, we first select lexical items that serve as group identifiers, attributes, and object terms, covering multiple stereotype domains. These lexical items are carefully curated from sociolinguistic literature and prior benchmarks to ensure coverage and relevance. Based on these lexical resources, we design a set of data templates containing placeholder slots for lexical insertion. Finally, we automatically generate a large number of evaluation instances by randomly sampling combinations of lexical items to populate the templates. This construction process ensures controlled manipulation of group membership and stereotype dimensions, enabling fine-grained measurement of model behavior under various social contexts.
3.3.1 Lexicon Resources Selection
We first select three widely studied stereotype domains: race, gender, and health. For each domain, we construct paired group identifiers and , representing advantaged and disadvantaged groups, respectively.
In the race domain, following prior studies Acerbi and Stubbersfield (2023), Bai et al. (2025) , we designated American as the advantaged group, while the disadvantaged groups were selected from four representative regions: African, Asian, Spanish-speaking, and Arab. In the gender domain, consistent with previous work Bai et al. (2025), females were treated as the advantaged group, while males and transgender individuals were considered disadvantaged groups. The health domain was further divided into four subdomains, where the advantaged groups included non-disabled, slim, young, and mentally healthy individuals, and the corresponding disadvantaged groups were disabled, overweight, aged, and individuals with mental illness. The specific group identifiers111https://www.behindthename.com/top/ used for each advantaged and disadvantaged category are summarized as follows in Table 1.
Next, we define positive attribute words () and negative attribute words () based on the 3 dimensions of the SCM: Competence, Sociability, and Morality. The initial set of attribute words was partially derived from prior studies Bai et al. (2025). Subsequently, we further invite five scholars with expertise in psychology to refine, supplement, and evaluate the attribute word lists. The final selection of attribute words is presented in Table 2.
| Dimension | Polarity | Attribute Words |
| Competence | (High Competence) | Proficient, Expertise, Efficient, Strategic, Resourceful, Innovative, Precise, Adaptable, Analytical, Competent, Insightful, Decisive, Masterful, Astute, Pioneering, Resilient, Impactful. |
| (Low Competence) | Incompetent, Inept, Unskilled, Weak, Deficient, Incapable, Ineffective, Powerless, Helpless, Feeble, Unqualified, Inadequate, Unfit, Untrained, Substandard, Unproficient, Lacking, Fragile, Mediocre, Undependable. | |
| Sociability | (High Sociability) | Outgoing, Sociable, Charismatic, Talkative, Approachable, Gregarious, Expressive, Enthusiastic, Collaborative, Convivial, Charming, Networked, Warm, Affable, Diplomatic, Engaging, People-oriented, Extroverted, Vivacious |
| (Low Sociability) | Reserved, Introverted, Quiet, Independent, Self-contained, Solitary, Contemplative, Private, Aloof, Detached, Reticent, Withdrawn, Unassuming, Pensive, Reclusive, Disengaged, Selective, Non-expressive, Insular | |
| Morality | (High Morality) | Principled, Ethical, Integrity-driven, Conscientious, Just, Accountable, Honorable, Scrupulous, Upright, Impartial, Dutiful, Righteous, Incorruptible, Law-abiding, Truthful, Reliable, Self-disciplined, Respectful, Steadfast, Dependable |
| (Low Morality) | Unprincipled, Unethical, Dishonest, Deceptive, Unaccountable, Corrupt, Unreliable, Duplicitous, Hypocritical, Negligent, Unscrupulous, Fraudulent, Deceitful, Manipulative, Unjust, Biased, Lawless, Self-serving, Exploitative, Opportunistic |
In addition to group identifiers and attribute words, we also constructed a set of neutral object words () to serve as fillers in the data templates. These neutral words were manually curated to represent inanimate and content-neutral entities that are not directly associated with any social group or stereotype. The words are presented in Table 3.
| Category | Words |
| Furniture | Table, Chair, Shelf |
| Vessel | Bottle, Plate, Cup, Box, Bag, Container |
| Tool | Pen, Key, Map, Coin, Wire, Pipe, Tool |
| Structure | Bridge, Window, Door, Frame, Fence |
| Nature | Road, Cloud, Stone, Hill, Path |
| Object | Book, Sheet, Lamp, Clock |
3.3.2 Data Template Design
For each task WABT and AAT, we design 3 sets of data templates to minimize potential unintended biases that may be introduced by specific prompt formulations. The data templates for each task are presented below.
For WABT task, we use color coding to distinguish different categories of words for clarity in the data templates. Specifically, green tags indicate either the advantaged group identifiers () or positive attribute words (), while red tags indicate either the disadvantaged group identifiers () or negative attribute words ().
In addition, blue tags are used to represent neutral object words () as well as group identity placeholders (), which may refer to either advantaged or disadvantaged social groups depending on the context for AAT task.
3.3.3 Data Generation
Finally, we perform automated construction by randomly sampling word combinations from the lexicon resources and inserting them into the data templates to generate the complete dataset.
Specifically, for the WABT task, we construct 10 paired combinations of and (e.g., African vs. American, Asian vs. American, etc.). For each combination, we randomly sample one pair of and group identifiers, and subsequently sample 5 and 5 attribute words from the lexicons corresponding to the three stereotype content dimensions. This sampling procedure is repeated 50 times for each combination. The sampled items are then combined with 3 data templates, resulting in a total of 4,500 instances. For the AAT task, we randomly sample 500 combinations of group identifiers and neutral nouns , and combine them with 3 data templates, resulting in a total of 1,500 instances. Figure 2 presents the distribution of the generated data.
4 Experiments and Results
4.1 Evaluated Models
We conduct evaluations on 8 mainstream open-source and closed-source LLMs, including LlaMa-2-70B-Chat Touvron et al. (2023), LlaMa-3-70B-Instruct Grattafiori et al. (2024), DeepSeek-V3 Liu et al. (2024) , DeepSeek-R1 Guo et al. (2025), GPT-4o Hurst et al. (2024), GPT-4-turbo, Claude-3.7-sonnet, and Gemini-2.5-pro.
4.2 Evaluation results on WABT
We input 4,500 data into 8 LLMs and obtain their respective responses. For each data, we record the number of valid responses returned by the models. For each valid response, we further compute the frequency counts of 4 specific combinations: , , , and . Based on these counts, we calculate the bias score along the 3 stereotype dimensions—Competence, Sociability, and Morality, following our predefined computational formulas.
After computing the bias scores for all data, we calculate the average bias score for each model along each stereotype dimension. We then conduct one-sample t-tests to assess whether the mean bias scores significantly deviated from 0. The results include the number of valid responses (), mean bias score (Mean), standard deviation of the bias scores (Std), t-statistic (), and significance level () for each model across the three dimensions, as summarized in the table. In general, larger -values indicate stronger bias tendencies, while smaller -values provide greater statistical confidence in the existence of such biases. The detailed experimental results are presented in Table 4.
Figure 3 provides a visual illustration of the bias score distributions across various social groups and stereotype dimensions for 4 LLMs with more than 1,000 valid responses.
Notably, distinct patterns emerge across models and dimensions. For example, some models exhibit pronounced negative biases in the Morality dimension toward specific groups (e.g., Disability or Overweight), whereas others display relatively neutral or even slightly positive bias scores.
Models Dimension Mean Std Claude-3.7-sonnet Competence 1470 -0.006 0.933 -0.263 0.792 Sociability 1482 0.415 0.830 19.273 .001 Morality 1465 -0.012 0.966 -0.493 0.622 DeepSeek-R1 Competence 178 -0.246 0.891 -3.672 .001 Sociability 183 0.320 0.787 5.492 .001 Morality 129 0.085 0.915 1.056 0.293 DeepSeek-V3 Competence 42 0.054 0.985 0.354 0.725 Sociability 17 0.180 0.945 0.763 0.456 Morality 23 0.130 0.991 0.617 0.544 Gemini-2.5-pro Competence 1351 -0.057 0.923 -2.252 0.025 Sociability 1398 0.367 0.792 17.329 .001 Morality 1352 -0.012 0.931 -0.489 0.625 GPT-4o Competence 382 -0.067 0.923 -1.423 0.156 Sociability 400 0.275 0.885 6.203 .001 Morality 419 0.176 0.956 3.770 .001 GPT-4-turbo Competence 1256 -0.063 0.953 -2.345 0.019 Sociability 1205 0.341 0.835 14.150 .001 Morality 1160 0.072 0.976 2.512 0.012 LlaMa-2-70B-Chat Competence 99 -0.039 0.698 -0.558 0.578 Sociability 113 0.256 0.657 4.124 .001 Morality 78 0.128 0.798 1.408 0.163 LlaMa-3-70B-Instruct Competence 1156 0.098 0.901 3.681 .001 Sociability 1130 0.246 0.854 9.682 .001 Morality 1095 0.065 0.920 2.337 0.020
4.3 Evaluation Results on AAT
We input 1,500 data into 8 LLMs and obtain their respective responses. For each data, we record the number of valid responses returned by the models. For each valid response, we further analyze the emotional framing chosen by the model—specifically, whether the response aligns more closely with a comedic or tragic interpretation.
Notably, a substantial portion of responses appear ambiguous or equivocal, indicating that the model does not make a clear choice between comedy and tragedy. We categorize such responses as Neutrality. We then compute the proportion of responses labeled as comedy, tragedy, and neutrality separately for cases where the social entity belongs to either or . The results are presented in Table 5.
Figure 4 presents the distribution of emotional framings across different social groups for each LLM, where green denotes the proportion of Comedy, red denotes Tragedy, and blue denotes Neutrality. The y-axis represents the percentage of each emotional category, with the total summing to 100% for each group.
The results reveal substantial variation in emotional framing across different social groups. Certain groups, such as Disability, Overweight, and Mental illness, are consistently associated with higher proportions of tragedy across multiple models, indicating a potential bias toward negatively valenced portrayals. In contrast, groups such as Asian, Youth, and American are more frequently linked with comedy or neutrality, suggesting relatively less stereotypical or emotionally charged representations.
Moreover, some models demonstrate particularly polarized patterns. For instance, DeepSeek-R1 and DeepSeek-V3 show overwhelmingly tragic framings across almost all groups, while LLaMa-2-70B-Chat produces a predominance of neutral responses, especially for marginalized identities.
Models Comedy Tragedy Neutrality Comedy Tragedy Neutrality Claude-3.7-sonnet 0.777 0.048 0.175 0.780 0.047 0.173 DeepSeek-R1 0.265 0.735 0.000 0.212 0.788 0.000 DeepSeek-V3 0.267 0.725 0.008 0.279 0.708 0.013 Gemini-2.5-pro 0.340 0.628 0.032 0.352 0.624 0.024 GPT-4o 0.267 0.725 0.008 0.279 0.708 0.013 GPT-4-turbo 0.532 0.465 0.003 0.452 0.547 0.001 LLaMa-2-70B-Chat 0.368 0.025 0.607 0.372 0.027 0.601 LLaMa-3-70B-Instruct 0.567 0.388 0.045 0.574 0.388 0.038
5 In-Depth Analysis of MIST
5.1 Consistently Observed: Pervasive Positive Bias in Sociability
One of the most prominent and consistent findings across the 8 LLMs evaluated in the WABT task is the widespread presence of positive bias in the Sociability dimension. With the exception of Claude-3.7-sonnet, whose bias score mean is close to 0, the majority of models exhibit a statistically significant positive Sociability bias (mean 0, p .001). Notably, Gemini-2.5-pro demonstrates the highest average bias score in Sociability (0.367) among all models, accompanied by a relatively low standard deviation (0.792), indicating a consistent tendency to attribute higher Sociability traits to a wide range of social groups.
Figure 3 provides further visual confirmation of this pattern. For models such as GPT-4-turbo, Gemini-2.5-pro, and LLaMa-3-70B-Instruct, the green lines representing the Sociability dimension extend outward across a broad spectrum of groups, including “Asian”, “African”, “Arab”, “Male”, and “Disabled”. This cross-model and cross-group consistency suggests a systematic inclination in the model behavior to portray entities as less sociable or friendly.
Such a tendency may stem from inherent biases in the training data, such as a preference for positive interpersonal interactions or idealized personality traits. Alternatively, it may reflect an inductive prior embedded in the model’s design, aimed at generating responses perceived as helpful, cooperative, or socially appropriate.
5.2 Multidimensional Complexity: Divergent Bias Patterns Across Dimensions
Bias in LLMs often exhibits considerable variation in direction, magnitude, and statistical significance across different stereotype dimensions, sometimes revealing independent or even opposing patterns. For example, on the Sociability dimension, DeepSeek-R1 demonstrates a strong and statistically significant positive bias, with an average bias score of 0.320 (p .001), indicating a consistent tendency to attribute higher Sociability traits to advantaged social targets. In contrast, the model exhibits a pronounced and significant negative bias on the Competence dimension (mean –0.246, p .001), suggesting a tendency to overestimate Competence in certain groups. Meanwhile, on the Morality dimension, the model yields an average score of 0.085 with no statistical significance (p 0.293), reflecting no consistent bias in that dimension. This striking contrast demonstrates that bias is not a monolithic structure, but rather a phenomenon that is independently manifested across different dimensions. Different attributes may exhibit independent, or even opposing, patterns of association.
Further evidence of this complexity can be found in other models. GPT-4o exhibits a significant positive bias in Sociability (mean 0.275, p .001) and also in Morality (mean 0.176, p .001), while its bias in Competence (mean –0.067) is statistically non-significant (p 0.156), indicating a relatively neutral stance in that dimension.
These cases highlight the heterogeneity of bias across models and dimensions, reinforcing the notion that bias in LLMs is inherently multi-faceted and dimension-specific, rather than uniformly expressed or aligned in a single direction.
5.3 Emergent Neutral Responses: Unexpected Patterns in Affective Attribution
In the AAT task design, the models are required to make a binary attribution decision between “Comedy” and “Tragedy”. However, in practice, several models spontaneously generate a proportion of outputs labeled as “Neutrality”, which are not pre-specified in the response options. This phenomenon indicates that some models exhibit attributional avoidance or uncertainty under certain social contexts.
As shown in Table 5, models produce near-zero proportions of “Neutrality” such as Claude-3.7-sonnet, LLaMa-3-70B-Instruct, and LLaMa-2-70B-Chat. Notably, LLaMa-2-70B-Chat yield the highest rate of neutral attributions, with 60.7% for groups and 60.1% for groups, making it the most concentrated model in terms of neutral responses. In contrast, both Claude-3.7-sonnet and LLaMa-3-70B-Instruct maintain relatively lower neutral rates (ranging from 4% to 17% across both and groups), but still demonstrated certain stable neutral tendencies under specific subgroup.
This neutral output phenomenon may reflect 2 potential mechanisms: (1) when models face certain sensitive social group identifiers, internal stereotype conflicts may lead to indecisive attribution behaviors, manifesting as ambiguous or uncertain attributional avoidance; (2) alternatively, some models may have been influenced by safety-oriented alignment optimization during training, causing them to proactively avoid emotionally sensitive outputs and instead favor neutralized responses as part of a “safety regulation avoidance mechanism”.
5.4 Asymmetry: Divergent Patterns of Implicit Bias
We first observe that across all evaluated language models, there is no simultaneous elevation in both FAR and UAR. This fundamental lack of simultaneous elevation demonstrates that the models’ affective attribution bias does not conform to a perfectly dual-peak distribution.
Further analysis reveals that multiple models exhibit relatively high bias levels on either the FAR or the UAR metric individually, rather than simultaneously showing high values on both metrics. For example, Claude-3.7-sonnet and LLaMa-3-70B-Instruct show higher scores on the FAR metric, reaching 77.7% and 56.7% respectively, indicating a stronger tendency to assign favorable attributions to advantaged groups, primarily driven by the amplification of positive affective associations. In contrast, the DeepSeek-R1 model achieves 78.8% on the UAR metric. This high score reflects a stronger tendency to assign unfavorable attributions to disadvantaged groups, predominantly through the amplification of negative affective associations.
This finding indicates that although all models exhibit group-level attribution bias at the global level, the underlying mechanisms through which such biases are expressed are not entirely consistent across models. Instead, models demonstrate divergent patterns in both the direction and magnitude of attribution. Some models predominantly exhibit “positive amplification for advantaged groups,” whereas others display “negative amplification for disadvantaged groups.”
6 Conclusion
In this paper, we designed a framework to evaluate implicit social biases in large language models by framing them as failures of Theory of Mind (ToM). We identified issues in existing evaluation methods, which often rely on direct-query tasks susceptible to social desirability effects and fail to capture the multi-dimensional nature of stereotypes. To address these, our method included two main strategies: (1) reconceptualizing bias through the multi-dimensional Stereotype Content Model (SCM), and (2) developing the Word Association Bias Test (WABT) and the Affective Attribution Test (AAT) as indirect tasks to elicit latent stereotypes. These strategies enabled our framework to probe for biases along distinct psychological dimensions and bypass the models’ explicit bias-avoidance mechanisms, improving the detection of subtle, structural stereotype patterns.