### v3

# MIST: Multi-dimensional Implicit BiaS Evaluation of LLMs for Theory of Mind

Abstract
Theory of Mind (ToM) in Large Language Models (LLMs) refers to the model’s ability to infer the mental states of others, with failures in this ability often manifesting as systemic implicit biases. Assessing this challenge is difficult, as traditional direct inquiry methods are often met with refusal to answer and fail to capture its subtle and multidimensional nature. Therefore, we propose MIST, which reconceptualizes the content model of stereotypes into multidimensional failures of ToM, specifically in the domains of competence, sociability, and morality. The framework introduces two indirect tasks. The Word Association Bias Test (WABT) assesses implicit lexical associations, while the Affective Attribution Test (AAT) measures implicit emotional tendencies, aiming to uncover latent stereotypes without triggering model avoidance. Through extensive experimentation on eight state-of-the-art LLMs, our framework demonstrates the ability to reveal complex bias structures and improved robustness. All data and code will be released.
WARNING: This paper contains content that may be offensive and disturbing in nature.
Keywords: Theory of Mind; Implicit Bias in Large Language Models; Stereotype Content Model
Introduction
As Large Language Models (LLMs) display more sophisticated reasoning, whether they possess a form of Theory of Mind (ToM) has become a central question [32]. ToM is the capacity to attribute and infer others’ intentions, beliefs, and mental states, and it requires inferences grounded in individualized, context-sensitive evidence. Under contemporary large-corpus pretraining, LLMs internalize group-level statistical regularities [5]. When these regularities are used to simplify social perception, they become imperfect generalizations in the form of stereotypes. If such group representations are inappropriately applied to individual-level inference, substituting group traits for person-specific evidence, bias arises and leads to systemic failures in ToM as shown in Figure 1. Therefore, identifying and quantifying bias is a prerequisite for evaluating and improving the ToM capacity of LLMs.
However, most existing studies [46], [10] employ uni-dimensional diagnostic tasks to evaluate such biases. Typically, these studies directly query the model to examine its associations with sensitive group-related attributes, such as gender, race, or occupation. While these approaches provide valuable empirical insights, prior studies [35, 40] have demonstrated that LLMs are also prone to social desirability effects in their responses. This susceptibility limits their ability to detect more subtle and cognitively plausible forms of bias that may surface during uncontrolled reasoning. Moreover, existing works [28, 23, 39, 36] have yet to adequately account for the multi-dimensional and relational nature of human social perception, which frequently involves the interplay of multiple psychological dimensions. This methodological gap results in a critical blind spot, potentially leading to an underestimation of how LLMs can perpetuate nuanced and socially corrosive stereotypes.
To address this research gap, we leverage the Stereotype Content Model (SCM), a widely employed framework in social psychology that characterizes stereotypes along three core dimensions: Competence, Sociability and Morality [12, 21]. This model provides a cognitively grounded multi-dimensional analytical framework for evaluating LLMs from a ToM perspective. Rather than directly probing for bias, we design indirect evaluation tasks Word Association Bias Test (WABT) and Affective Attribution Test (AAT). Specifically, WABT measures associative biases by having the model pair attribute words with social groups, while AAT measures affective biases by having it attribute an emotional valence to generated scenarios involving those groups. By framing the evaluations as objective lexical association or subjective affective judgment tasks, rather than direct inquiries about social beliefs, the methodology avoids triggering the model’s learned social desirability filters. Consequently, these tasks naturally prompt the model to generate social inferences without explicitly introducing the notion of bias, thereby allowing its latent stereotypical tendencies to surface unconsciously during the reasoning process.
The contributions of this paper are summarized as follows:
We provide a theoretical perspective that combines insights from ToM and SCM to reconceptualize implicit bias in LLMs as systematic failures in mental state modeling.
We propose a novel implicit bias evaluation framework that incorporates WABT and AAT tasks, which indirectly prompt the model to generate group-level inferences along the 3 SCM dimensions. This design minimizes the influence of explicit bias-avoidance mechanisms.
We conduct extensive empirical evaluations on multiple SOTA LLMs, uncovering new insights into the structural, subtle, and pervasive nature of their implicit social biases.
Evaluation Framework
Tasks Definition
We design two types of implicit bias evaluation tasks Word Association Bias Test (WABT) and Affective Attribution Test (AAT), to assess LLMs’ implicit biases and underlying stereotypical tendencies along the three bias and stereotype dimensions of Competence, Sociability, and Morality.
WABT. The WABT indirectly assesses LLMs’ implicit biases by examining their associative tendencies at the lexical level. These implicit biases and stereotypes are often reflected in the model’s inclination to associate specific groups with certain attributes or characteristics when processing group-related words. Specifically, for each bias dimension, a pair of target group identifiers and 10 attribute words (5 , 5 ) are provided to model. The model is required to associate each attribute word with one of the two target group identifiers. The model’s output is represented as pairs. refers to the positively framed target group (advantaged or normative group), and refers to the negatively framed target group (disadvantaged or marginalized group). Likewise, denotes positive or desirable attributes, while denotes negative or undesirable attributes. WABT examines whether LLMs reflect an understanding or misinterpretation of others’ mental states through lexical associations, linking to the ToM in social interaction and group cognition reasoning.
AAT. The AAT is designed to evaluate LLMs’ implicit biases and stereotypes by examining their affective associations toward social group identifiers. The task is adapted from the Affect Misattribution Procedure in cognitive psychology, which infers implicit attitudes based on affective priming effects. Specifically, for each social group dimension, is first prompted in each trial to generate a descriptive sentence that includes a neutral word alongside the target group identifier which is a combination of and . Subsequently, the model is required to categorize the generated sentence, based on its initial affective response, into one of two categories: Comedy (positive valence) or Tragedy (negative valence). The output of is recorded as a categorical label, reflecting the affective association activated toward the target group. AAT examines whether LLMs can identify and infer emotional responses based on social group membership, reflecting its social cognition reasoning abilities within the context of ToM.
Evaluation Metrics
For each task, we employ specific evaluation metrics to rigorously quantify the extent of implicit bias exhibited by LLMs.
WABT. To quantify the implicit association bias in each test, we adopt a commonly used lexical association bias scoring method. The bias score is computed as follows:
where denotes the number of times the model assigns an attribute word from to the target group (i.e., the number of pairs in the model’s output), and similarly for the other terms. The resulting bias score ranges from (completely reversed bias) to (completely consistent bias), with indicating no observable bias.
AAT. We focus on model’s affective attribution tendencies along two target-specific directions: (a) When target group identifier belongs to (advantaged group) and the model classifies it as comedy (positive valence), it is counted as a favorable attribution. (b) When target group identifier belongs to (disadvantaged group) and model classifies it as tragedy (negative valence), it is counted as an unfavorable attribution.
After multiple rounds of testing, we obtain the number of favorable attributions, denoted as , and the number of unfavorable attributions, denoted as . Following the collection of the total number of favorable and unfavorable attributions, we further compute two normalized attribution rates: (a) Favorable Attribution Rate (FAR) and (b) Unfavorable Attribution Rate (UAR). The FAR as the proportion of favorable attributions among the total number of instances where the target group belongs to (advantaged group), defined as The UAR as the proportion of unfavorable attributions among the total number of instances where the target group belongs to (disadvantaged group), defined as Intuitively, higher values of FAR and UAR indicate stronger implicit biases and stereotypical tendencies in the model’s affective attribution behavior. Specifically, a high FAR suggests that the model disproportionately associates advantaged groups () with positive valence (Comedy), while a high UAR reflects a tendency to associate disadvantaged groups () with negative valence (Tragedy). Both patterns reveal systematic asymmetries in the model’s social reasoning that may reflect internalized societal stereotypes.
Data Construction
We further construct a synthetic dataset through a controlled generation pipeline. Specifically, we first select lexical items serving as group identifiers, attribute terms, and object terms, covering multiple stereotype domains. These lexical items are carefully curated from ToM and sociological literature, as well as existing bias evaluation benchmarks [31, 29], to ensure both coverage and relevance. Based on these lexical resources, we design a set of data templates with placeholder slots for lexical insertion. Finally, a large number of evaluation instances are automatically generated by randomly sampling combinations of lexical items to populate the templates. This construction process enables controlled manipulation of group membership and stereotype dimensions, allowing for fine-grained measurement of model behavior across diverse social contexts.
Lexicon Resources Selection. We first select three widely studied bias domains: race, gender, and health. For each domain, we construct paired group identifiers and (cf. Table 1), representing advantaged and disadvantaged groups. In the race domain, following prior studies [1, 4], we designate American as the advantaged group, while the disadvantaged groups are selected from four representative regions: African, Asian, Spanish-speaking, and Arab. In the gender domain, consistent with previous work [4], females are treated as the advantaged group, while males and transgender individuals are considered disadvantaged groups. The health domain is further divided into four subdomains, where the advantaged groups included non-disabled, slim, young, and mentally healthy individuals, and the corresponding disadvantaged groups were disabled, overweight, aged, and individuals with mental illness. Next, we define positive attribute words () and negative attribute words () based on the 3 dimensions of the SCM. The initial set of attribute words was partially derived from prior studies [4]. Subsequently, we further invite five scholars with expertise in psychology to refine, supplement, and evaluate the attribute word lists (cf. Table 2). In addition to group identifiers and attribute words, we also construct a set of neutral object words () to serve as fillers in the data templates. These neutral words are manually curated to represent inanimate and content-neutral entities that are not directly associated with any social group or stereotype (cf. Table 3).
| Domain | Subdomain | Group | Category | Items |
| Race | - | American | Ethan, Olivia, Jackson, Ava, Liam | |
| African | Kwame, Amina, Zuberi, Thabo, Adama | |||
| Asian | Hiroshi, Jiahao, Aditi, Siti, Minh | |||
| Spanish-speaking | Martín, Valentina, Diego, Camila, Felipe | |||
| Arab | Omar, Layla, Rami, Nadia, Samir | |||
| Gender | - | Female | Woman, Girl, Lady, Feminine, Womanhood | |
| Male | Man, Boy, Gentleman, Masculine, Manhood | |||
| Transgender | Trans, Gender-nonconforming, Genderqueer, Non-binary, Gender-fluid | |||
| Health | Disability | - | Able, Fitness, Wellness | |
| - | Disable, Impairment, Handicap | |||
| Weight | - | Slim, Skinny, Thin | ||
| - | Obesity, Heavy, Fat | |||
| Age | - | Young, Youth, Junior | ||
| - | Old, Aged, Seniors | |||
| Mind | - | Sanity, Emotional Stability, Sane | ||
| - | Depression, Schizophrenia, Bipolar Disorder |
| Dimension | Polarity | Attribute Words |
| Competence | (High Competence) | Proficient, Expertise, Efficient, Strategic, Resourceful, Innovative, Precise, Adaptable, Analytical, Competent, Insightful, Decisive, Masterful, Astute, Pioneering, Resilient, Impactful. |
| (Low Competence) | Incompetent, Inept, Unskilled, Weak, Deficient, Incapable, Ineffective, Powerless, Helpless, Feeble, Unqualified, Inadequate, Unfit, Untrained, Substandard, Unproficient, Lacking, Fragile, Mediocre, Undependable. | |
| Sociability | (High Sociability) | Outgoing, Sociable, Charismatic, Talkative, Approachable, Gregarious, Expressive, Enthusiastic, Collaborative, Convivial, Charming, Networked, Warm, Affable, Diplomatic, Engaging, People-oriented, Extroverted, Vivacious |
| (Low Sociability) | Reserved, Introverted, Quiet, Independent, Self-contained, Solitary, Contemplative, Private, Aloof, Detached, Reticent, Withdrawn, Unassuming, Pensive, Reclusive, Disengaged, Selective, Non-expressive, Insular | |
| Morality | (High Morality) | Principled, Ethical, Integrity-driven, Conscientious, Accountable, Honorable, Scrupulous, Upright, Impartial, Dutiful, Righteous, Incorruptible, Law-abiding, Truthful, Reliable, Self-disciplined, Respectful, Steadfast, Dependable |
| (Low Morality) | Unprincipled, Unethical, Dishonest, Deceptive, Unaccountable, Corrupt, Unreliable, Duplicitous, Hypocritical, Negligent, Unscrupulous, Fraudulent, Deceitful, Manipulative, Unjust, Lawless, Self-serving, Exploitative, Opportunistic |
| Category | Words |
| Furniture | Table, Chair, Shelf |
| Vessel | Bottle, Plate, Cup, Box, Bag, Container |
| Tool | Pen, Key, Map, Coin, Wire, Pipe, Tool |
| Structure | Bridge, Window, Door, Frame, Fence |
| Nature | Road, Cloud, Stone, Hill, Path |
| Object | Book, Sheet, Lamp, Clock |
Data Template Design. For each task, we design 3 sets of data templates to minimize potential unintended biases that may be introduced by specific prompt formulations.
Data Generation. Finally, we perform automated construction by randomly sampling word combinations from the lexicon resources and inserting them into data templates to generate complete dataset. Specifically, for the WABT, we construct 10 paired combinations of and (e.g., African vs. American, Asian vs. American, etc.). For each combination, we randomly sample one pair of and group identifiers, and subsequently sample 5 and 5 attribute words from the lexicons corresponding to the three stereotype content dimensions. This sampling procedure is repeated 50 times for each combination. The sampled items are then combined with 3 data templates, resulting in a total of 4,500 instances. For the AAT, we randomly sample 500 combinations of group identifiers and neutral nouns , and combine them with 3 data templates, resulting in a total of 1,500 instances.
Experiment Results
Evaluated Models
Evaluation results on WABT
We input each instance into LLMs and obtain their respective responses. For each data, we record the number of valid responses returned by the models. For each valid response, we further compute the frequency counts of 4 specific combinations: , , , and . Based on these counts, we calculate the bias score, following our predefined computational formulas.
After computing the bias scores for all data, we calculate the average bias score for each model along each stereotype dimension. We then conduct one-sample t-tests to assess whether the mean bias scores significantly deviated from 0. The results include the number of valid responses (), mean bias score (Mean), standard deviation of the bias scores (Std), t-statistic (), and significance level () for each model across the three dimensions, as summarized in the table. In general, larger -values indicate stronger bias tendencies, while smaller -values provide greater statistical confidence in the existence of such biases. The detailed results are presented in Table 4.
Notably, distinct patterns emerge across models and dimensions. For example, some models exhibit pronounced negative biases in the Morality dimension toward specific groups (e.g., Disability or Overweight), whereas others display relatively neutral or even slightly positive bias scores.
| Models | Dim. | Mean | Std | |||
| Claude-3.7-sonnet | Com. | 1470 | -0.006 | 0.933 | -0.263 | 0.792 |
| Soc. | 1482 | 0.415 | 0.830 | 19.273 | .001 | |
| Mor. | 1465 | -0.012 | 0.966 | -0.493 | 0.622 | |
| DeepSeek-R1 | Com. | 178 | -0.246 | 0.891 | -3.672 | .001 |
| Soc. | 183 | 0.320 | 0.787 | 5.492 | .001 | |
| Mor. | 129 | 0.085 | 0.915 | 1.056 | 0.293 | |
| DeepSeek-V3 | Com. | 42 | 0.054 | 0.985 | 0.354 | 0.725 |
| Soc. | 17 | 0.180 | 0.945 | 0.763 | 0.456 | |
| Mor. | 23 | 0.130 | 0.991 | 0.617 | 0.544 | |
| Gemini-2.5-pro | Com. | 1351 | -0.057 | 0.923 | -2.252 | 0.025 |
| Soc. | 1398 | 0.367 | 0.792 | 17.329 | .001 | |
| Mor. | 1352 | -0.012 | 0.931 | -0.489 | 0.625 | |
| GPT-4o | Com. | 382 | -0.067 | 0.923 | -1.423 | 0.156 |
| Soc. | 400 | 0.275 | 0.885 | 6.203 | .001 | |
| Mor. | 419 | 0.176 | 0.956 | 3.770 | .001 | |
| GPT-4-turbo | Com. | 1256 | -0.063 | 0.953 | -2.345 | 0.019 |
| Soc. | 1205 | 0.341 | 0.835 | 14.150 | .001 | |
| Mor. | 1160 | 0.072 | 0.976 | 2.512 | 0.012 | |
| LlaMa-2-70B-Chat | Com. | 99 | -0.039 | 0.698 | -0.558 | 0.578 |
| Soc. | 113 | 0.256 | 0.657 | 4.124 | .001 | |
| Mor. | 78 | 0.128 | 0.798 | 1.408 | 0.163 | |
| LlaMa-3-70B-Instruct | Com. | 1156 | 0.098 | 0.901 | 3.681 | .001 |
| Soc. | 1130 | 0.246 | 0.854 | 9.682 | .001 | |
| Mor. | 1095 | 0.065 | 0.920 | 2.337 | 0.020 |
Evaluation Results on AAT
We input each data into 8 LLMs and obtain their respective responses. For each instance, we record the number of valid responses returned by the models. For each valid response, we further analyze the emotional framing chosen by the model—specifically, whether the response aligns more closely with a comedic or tragic interpretation.
Notably, a substantial portion of responses appear ambiguous or equivocal, indicating that the model does not make a clear choice between comedy and tragedy. We categorize such responses as Neutrality. We then compute the proportion of responses labeled as comedy, tragedy, and neutrality separately for cases where the social entity belongs to either or . The results are presented in Table 5. The results reveal substantial variation in emotional framing across different social groups. Certain groups, such as Disability, Overweight, and Mental illness, are consistently associated with higher proportions of tragedy across multiple models, indicating a potential bias toward negatively valenced portrayals. In contrast, groups such as Asian, Youth, and American are more frequently linked with comedy or neutrality, suggesting relatively less stereotypical or emotionally charged representations. Moreover, some models demonstrate particularly polarized patterns. For instance, DeepSeek-R1 and DeepSeek-V3 show overwhelmingly tragic framings across almost all groups, while LLaMa-2-70B-Chat produces a predominance of neutral responses, especially for marginalized identities.
| Models | ||||||
| C. | T. | N. | C. | T. | N. | |
| Claude-3.7-sonnet | 0.777 | 0.048 | 0.175 | 0.780 | 0.047 | 0.173 |
| DeepSeek-R1 | 0.265 | 0.735 | 0.000 | 0.212 | 0.788 | 0.000 |
| DeepSeek-V3 | 0.267 | 0.725 | 0.008 | 0.279 | 0.708 | 0.013 |
| Gemini-2.5-pro | 0.340 | 0.628 | 0.032 | 0.352 | 0.624 | 0.024 |
| GPT-4o | 0.267 | 0.725 | 0.008 | 0.279 | 0.708 | 0.013 |
| GPT-4-turbo | 0.532 | 0.465 | 0.003 | 0.452 | 0.547 | 0.001 |
| LLaMa-2-70B-Chat | 0.368 | 0.025 | 0.607 | 0.372 | 0.027 | 0.601 |
| LLaMa-3-70B-Instruct | 0.567 | 0.388 | 0.045 | 0.574 | 0.388 | 0.038 |
In-Depth Analysis of MIST
Finding-1: A consistently observed pervasive positive bias in sociability. One of the most prominent and consistent findings in the WABT is the widespread presence of positive bias in the Sociability dimension. With the exception of Claude-3.7-sonnet, whose bias score mean is close to 0, the majority of models exhibit a statistically significant positive Sociability bias (mean 0, p .001). Notably, Gemini-2.5-pro demonstrates the highest average bias score in Sociability (0.367) among all models, accompanied by a relatively low standard deviation (0.792), indicating a consistent tendency to attribute higher Sociability traits to a wide range of social groups. Such a tendency may stem from inherent biases in the training data, such as a preference for positive interpersonal interactions or idealized personality traits. Alternatively, it may reflect an inductive prior embedded in the model’s design, aimed at generating responses perceived as helpful, cooperative, or socially appropriate.
Finding-2: Multidimensional complexity of divergent bias patterns across dimensions. Bias in LLMs varies substantially across stereotype dimensions in terms of direction, magnitude, and statistical significance, and may even exhibit opposing patterns. For example, DeepSeek-R1 shows a significant positive bias in Sociability (0.320, p .001), while exhibiting a significant negative bias in Competence (–0.246, p .001). In contrast, its bias in Morality is not statistically significant (0.085, p 0.293). This contrast indicates that bias patterns are not synchronized across dimensions. Similar heterogeneity is observed in other models. GPT-4o demonstrates significant positive bias in both Sociability and Morality, while its bias in Competence is statistically non-significant, suggesting a relatively neutral stance in that dimension. Overall, these results show that bias in LLMs is not a monolithic phenomenon, but rather a multi-faceted, dimension-specific structure.
Finding-3: Emergent neutral response patterns in affective attribution. In the AAT, models are instructed to make a binary affective attribution between Comedy and Tragedy. However, several models spontaneously generate a non-trivial proportion of Neutrality responses that are not specified in the predefined output space, indicating attributional avoidance or uncertainty under certain social contexts. As shown in Table 5, substantial variation exists across models in the prevalence of neutral outputs. LLaMa-2-70B-Chat exhibits the highest concentration of neutral attributions, reaching 60.7% for groups and 60.1% for groups. In contrast, Claude-3.7-sonnet and LLaMa-3-70B-Instruct maintain comparatively low neutral rates (approximately 4%–17%), though they still display stable neutral tendencies within specific subgroups. This emergent neutrality may reflect two underlying mechanisms: (1) exposure to sensitive social group identifiers induces internal stereotype conflicts, leading to indecisive or avoidance-oriented attribution behavior; and (2) safety-oriented alignment objectives encourage models to proactively avoid emotionally sensitive judgments, favoring neutralized responses as a form of implicit safety regulation.
Finding-4: Implicit bias exhibits asymmetric and divergent patterns. Across all evaluated LLMs, FAR and UAR do not increase simultaneously, indicating that affective attribution bias does not follow an idealized dual-peak distribution. Instead, most models exhibit elevated bias on only one of the two metrics. For instance, Claude-3.7-sonnet and LLaMa-3-70B-Instruct achieve higher FAR scores of 77.7% and 56.7%, respectively, suggesting a stronger tendency to assign favorable affective attributions to advantaged groups. In contrast, DeepSeek-R1 reaches 78.8% on the UAR metric, reflecting a stronger tendency to assign unfavorable affective attributions to disadvantaged groups. These results indicate that although all models display group-level attribution bias at a global level, the mechanisms through which such biases are expressed differ across models. Some models are primarily characterized by positive amplification toward advantaged groups, whereas others exhibit negative amplification toward disadvantaged groups.
Related Work
ToM in LLMs. Recent advancements in LLM reasoning [44, 43, 17, 16] have spurred debate on their potential for an emergent ToM. Evaluating this capacity is now a pivotal goal for human-centered AI [19, 22, 26, 41, 7, 25]. However, scholarly assessments diverge. Some critics argue that high performance stems from flawed benchmarks or superficial statistical patterns rather than genuine reasoning [42, 33], while other research shows ToM can be robustly trained to generalize [27]. A critical failure arises when LLMs misapply learned societal stereotypes to individuals.
Stereotype Content Model. SCM explains stereotypes along Warmth and Competence [11], with Warmth refined into Sociability and Morality [21], highlighting ambivalent prejudice [9, 13, 6]. Recent work shows that LLMs reproduce SCM-consistent patterns: despite generally positive tone, group descriptions align with SCM dimensions [20, 34], defaulting to white, healthy, middle-aged male representations while exhibiting semantic shifts for other groups [4, 37]. This work [30] analyses further indicate that Warmth and Competence remain dominant evaluative axes and that bias extends beyond binary labels, with implications for education and hiring [2, 45].
Conclusion
In this paper, we design a framework MIST to evaluate implicit social biases in LLMs by framing them as failures of ToM. We identify issues in existing evaluation methods, which often rely on direct-query tasks susceptible to social desirability effects and fail to capture the multi-dimensional nature of stereotypes. To address these, our method includes two main strategies: (1) reconceptualizing bias through the multi-dimensional Stereotype Content Model, and (2) developing the Word Association Bias Test and the Affective Attribution Test as indirect tasks to elicit latent biases. These strategies enable our framework to probe for biases along distinct psychological dimensions and bypass the models’ explicit bias-avoidance mechanisms, improving the detection of subtle, structural stereotype patterns.